# Also at ]Alpine Quantum Technologies GmbH, Innsbruck, Austria

### v1

@@ II.2 Ideal tomography for the dynamical error map
$\mathcal{E}_{t}(\rho_{0})=\sum_{n=1}^{d^{2}}{K}_{n}(t)\rho_{\rm id}(t)\,{K}^{\dagger}_{n}(t),\hskip 8.61108pt\sum_{n=1}^{d^{2}}{K}_{n}^{\dagger}(t){K}_{n}(t)=\mathbb{1}_{d}.$ (26)
@@
$p_{s,i,\mu}={\rm tr}\left(D_{s\mu}\chi(t_{i})\right),\hskip 4.30554pt[D_{s\mu}]_{\beta\alpha}={\rm tr}\left(M_{\mu}E_{\alpha}\rho_{0,s}E_{\beta}^{\dagger}\right).$ (29)
Here, the set of $d^{2}\times d^{2}$ matrices $\{D_{s\mu}:\mu\in\mathbb{M}_{f},s\in\mathbb{S}_{0}\}$ contains all of the information about the state preparation and readout. We remark that the estimated $\chi(t_{i})$ is the process matrix for the complete evolution, and not that for the error channel that takes place after the ideal unitary gate $U_{g}(t_{i})$. Therefore, it must be diagonalized to obtain the corresponding error Kraus operators, which are then transformed to express the full dynamical map at $t_{i}$ as the ideal gate $U_{g}(t_{i})=U_{\Omega}(t_{i})$ followed by the error in Eq. (26) with
$K_{n}(t_{i})=\sqrt{d_{n}(t_{i})}\sum_{\alpha=1}^{d^{2}}{v}_{\alpha n}(t_{i})U_{\Omega}(t_{i})E_{\alpha}U_{\Omega}^{\dagger}(t_{i}),$ (30)
@@
Remarkably, using the analytical solutions of the dressed-state master equation that depend on the filtered integrals in Eq. (18), we have found closed analytical expressions for the Kraus operators that provide a concise description of the effect colored noise has on the ideal unitary evolution $\forall t_{i}\in T$, which we then label again as $t$. To limit the complexity of the expressions, we start by disregarding the contributions of the last two filter functions in Eq. (24), whose influence can be vanishingly small in the Markovian limit. This amounts to setting $\Gamma_{2}(t)=\Delta_{2}(t)=0$ in Eq. (97), such that
@@
Here, $\gamma_{1}$, $\gamma_{2}$, and $\delta_{2}$ encode the effect of the dephasing noise on the qubit via its correlation function, as discussed in detail in Eqs. (91) and (94) in Appendix C. The second rate is always positive $\bar{\gamma}_{+}(t)=-\textstyle\frac{1}{2}\partial_{t}\log(1-\epsilon(t))>0\,\,\forall t$, which means it never contributes to the value of the integral. Clearly, when $\gamma_{2}(t)=\delta_{2}(t)=0$, one gets $\bar{\gamma}_{-}(t)=\bar{\gamma}_{+}(t)>0$, such that the non-Markovianity measure vanishes, signaling a CP-divisible dynamical map. Hence, including the additional filtered integrals in Eq. (24) is crucial to model non-Markovian noise effects in the gates. We depict the non-Markovianity measure in Fig. 3 for an Ornstein-Uhlenbeck process describing the phase noise. It is interesting to mention that $\mathcal{N}_{\rm CP}$ saturates at the same time for all the Rabi frequencies shown, and that it possesses a unique maximum at a certain Rabi frequency.
@@
$\begin{split}\Gamma_{1}(t)&\approx\frac{t}{T_{\rm 2,eff}},\hskip 17.22217ptT_{2,\rm eff}=\frac{2}{S(\Omega)},\\ \Delta_{1}(t)&\approx\delta\Omega\,t,\hskip 21.52771pt\delta\Omega=\Omega\frac{\tau_{\rm c}}{T_{2,\rm eff}}.\end{split}$ (37)
@@
Figure 4: Comparison of average channel infidelities, capturing the deviations of the full numerical solutions of the stochastic Langevin equations with the depolarizing channel (red), the Pauli-twirled channel (purple), the non-Clifford Kraus channel (blue), and the full non-Clifford and non-Markovian channel (orange). We average the simulated Langevin SDEs over $M_{\rm MC}=10^{5}$ noise trajectories, and average the channel infidelities by randomly drawing $N_{0}=5\times 10^{3}$ initial pure states. The simulations use an Ornstein-Uhlenbeck dephasing noise with $\tau_{\rm c}=20\pi/\Omega$, $c=1/(40\tau_{\rm c}^{3})$, a time step $\Delta t=0.05\tau\,{\rm c}$, and a Rabi frequency $\Omega=2\pi\times 20\,{\rm kHz}$, which are chosen for presentation purposes and are not connected to the experimental trapped-ion noise of the following section. The local minima appearing in the effective Kraus channel are due to the over-rotations induced by the non-Markovian effects that this approximation does not capture.
@@
II.3 Projection noise, maximum-likelihood, and analytical estimates of the gate error rate
@@
Figure 5: Reconstruction of the gate error $\overline{\epsilon}(t_{i})$ for noisy Rabi oscillations under an Ornstein-Uhlenbeck noise with $\tau_{c}=5\times 10^{-4}\,\rm{s}$ and $c=1.6\times 10^{9}\,\rm{s}^{-3}$. We include shot noise for a total number of measurements $N_{\rm shots}=2.4\cdot 10^{7}$, which allows us to approximate the tomographic probabilities $p_{s,i,\mu}$ by the corresponding relative frequencies $f_{s,i,\mu}$. The confidence region (purple shading) and the mean (purple markers) are estimated by distributing this data into $N_{r}=10^{3}$ repetitions of the maximum-likelihood reconstruction, leading to $N_{i}=1200$ shots per time step. The analytical predictions $\epsilon_{g,{\rm D}}$ in Eq. (43) (green), as well as the errors for our Markovian non-Clifford $\epsilon_{g,{\rm NC}}$ in Eq. (44) (blue) and non-Markovian $\epsilon_{g,{\rm NM}}$ in Eq. (45) (red) estimates, which depend on the Lorentzian noise PSD in Eq. (25). Inset: Absolute value of the difference between the ML reconstruction and the non-Clifford Kraus map (blue) or the non-Markovian map (red) infidelity.
@@
This expression captures better the oscillations of the gate error, as can be seen in the inset of Fig. 5. We note that the Pauli-twirled approximation $\mathcal{E}_{\rm PT}$ in Eq. (40) has the exact same fidelity but, being Pauli-like, can be used in large-scale simulations of Clifford circuits.

### v3

@@ II.2 Ideal tomography for the dynamical error map
$\mathcal{E}_{t}(\rho_{0})=\sum_{n=1}^{d^{2}}{K}^{\phantom{\dagger}}_{n}(t)\rho_{\rm id}(t)\,{K}^{\dagger}_{n}(t),\hskip 8.61108pt\sum_{n=1}^{d^{2}}{K}_{n}^{\dagger}(t){K}^{\phantom{\dagger}}_{n}(t)=\mathbb{1}_{d}.$ (26)
@@
$p_{s,i,\mu}={\rm tr}\left(D_{s\mu}\chi(t_{i})\right),\hskip 4.30554pt[D_{s\mu}]_{\beta\alpha}={\rm tr}\left(M_{\mu}E_{\alpha}^{\phantom{\dagger}}\rho_{0,s}E_{\beta}^{\dagger}\right).$ (29)
Here, the set of $d^{2}\times d^{2}$ matrices $\{D_{s\mu}:\mu\in\mathbb{M}_{f},s\in\mathbb{S}_{0}\}$ contains all of the information about the state preparation and readout. We remark that the estimated $\chi(t_{i})$ is the process matrix for the complete evolution, and not just that for the error channel taking place after the ideal unitary gate $U_{g}(t_{i})$. Therefore, it must be diagonalized to obtain the corresponding error Kraus operators, which are then transformed to express the full dynamical map at $t_{i}$ as the ideal gate $U_{g}(t_{i})=U_{\Omega}(t_{i})$ followed by the error in Eq. (26) with
$K_{n}(t_{i})=\sqrt{d_{n}(t_{i})}\sum_{\alpha=1}^{d^{2}}{v}_{\alpha n}(t_{i})U^{\phantom{\dagger}}_{\Omega}(t_{i})E_{\alpha}U_{\Omega}^{\dagger}(t_{i}),$ (30)
@@
Remarkably, using the analytical solutions of the dressed-state master equation that depend on the filtered integrals in Eq. (18), we have found closed analytical expressions for the Kraus operators that provide a concise description of the effect colored noise has on the ideal unitary evolution $\forall t_{i}\in T$, which we then label again as $t$. To limit the complexity of the expressions, we start by disregarding the contributions of the last two filter functions in Eq. (LABEL:eq:noN_tarkov_filter_functions), whose influence can be vanishingly small in the Markovian limit. This amounts to setting $\Gamma_{2}(t)=\Delta_{2}(t)=0$ in Eq. (17), such that
@@
Here, $\gamma_{1}$, $\gamma_{2}$, and $\delta_{2}$ encode the effect of the dephasing noise on the qubit via its correlation function, as discussed in detail in Eqs. (104) and (107) in Appendix C. The second rate is always positive $\bar{\gamma}_{+}(t)=-\textstyle\frac{1}{2}\partial_{t}\log(1-\epsilon(t))>0\,\,\forall t$, which means it never contributes to the value of the integral. Clearly, when $\gamma_{2}(t)=\delta_{2}(t)=0$, one gets $\bar{\gamma}_{-}(t)=\bar{\gamma}_{+}(t)>0$, such that the non-Markovianity measure vanishes, signaling a CP-divisible dynamical map. Hence, including the additional filtered integrals in Eq. (LABEL:eq:noN_tarkov_filter_functions) is crucial to model non-Markovian noise effects in the gates. We depict the non-Markovianity measure in Fig. 3 for an Ornstein-Uhlenbeck process describing the phase noise. It is interesting to mention that $\mathcal{N}_{\rm CP}$ saturates at the same time for all the Rabi frequencies shown, and that it possesses a unique maximum at a certain Rabi frequency.
@@
$\begin{split}\Gamma_{1}(t)&\approx\frac{t}{T_{\rm 2,eff}},\hskip 17.22217ptT_{2,\rm eff}=\frac{2}{S_{\tilde{\omega}}(\Omega)},\\ \Delta_{1}(t)&\approx\delta\Omega\,t,\hskip 21.52771pt\delta\Omega=\Omega\frac{\tau_{\rm c}}{T_{2,\rm eff}}.\end{split}$ (37)
@@
Figure 4: Comparison of average channel infidelities, capturing the deviations of the full numerical solutions of the stochastic Langevin equations with the depolarizing channel (red), the Pauli-twirled channel (purple), the non-Clifford Kraus channel (blue), and the full non-Clifford and non-Markovian channel (orange). We average the simulated Langevin SDEs over $M_{\rm MC}=10^{5}$ noise trajectories, and average the channel infidelities by randomly drawing $N_{0}=5\times 10^{3}$ initial pure states. The simulations use an Ornstein-Uhlenbeck dephasing noise with $\tau_{\rm c}=20\pi/\Omega$, $c=1/(40\tau_{\rm c}^{3})$, a time step $\Delta t=0.05\tau_{\rm c}$, and a Rabi frequency $\Omega=2\pi\times 20\,{\rm kHz}$, which are chosen for presentation purposes and are not connected to the experimental trapped-ion noise of the following section. The local minima appearing in the effective Kraus channel are due to the over-rotations induced by the non-Markovian effects that this approximation does not capture.
@@
II.3 Dynamical error map with amplitude noise
Up until now we have focused entirely on phase and frequency noise. However, real quantum systems are typically subject to various types of noise simultaneously. If the underlying noise operators commute and are mutually uncorrelated then the dynamical map of the net noise effect is simply the composition of the individual noise maps. Unfortunately, this is rarely the case, and the treatment of universal or multi-axis noise becomes more difficult in general. As we now show, it is possible to generalize our formalism to incorporate the effects of transversal noise, such that both amplitude and phase fluctuations in the drive affect the single-qubit gates. The drive amplitude fluctuations will be modeled by stochastic noise in the Rabi frequency [120] roughly as
$\Omega\to\Omega+\delta\tilde{\Omega}(t),\hskip 4.30554pt\delta\tilde{\Omega}(t)\simeq\frac{\Omega}{2}\frac{\delta\tilde{I}_{L}(t)}{I_{L,0}},$ (42)
where $\delta\tilde{\Omega}(t)$ and $\delta\tilde{I}_{L}(t)$ denote the stochastic fluctuations of the Rabi rate and the laser intensity, respectively, and $I_{L,0}$ denotes the ideal intensity of the driving field, where intensity is proportional to the square of the field amplitude.
The Hamiltonian governing the dynamics of the qubit under a driving with both phase and amplitude fluctuations can be written in a toggling frame, within the same approximations as in Eq. (6), as
$\tilde{H}(t)=\frac{1}{2}\left(\Omega+\delta\tilde{\Omega}(t)\right)\sigma_{\phi}-\frac{1}{2}\delta\tilde{\omega}(t)\sigma_{z}.$ (43)
Following similar steps as in our previous derivation, we can approximate the dynamics of the system using a cluster expansion of the Nakajima-Zwanzig equation. For simplicity we pick in the following the case $\phi=0$ such that the Rabi drive is around $\sigma_{x}$ again and we consider stochastic processes with zero mean. We also assume that the processes are individually correlated in time, but display no cross-correlations, i.e., $\mathbb{E}\{\delta\tilde{\omega}(t)\delta\tilde{\Omega}(t^{\prime})\}=0$, leaving the more-involved case of arbitrary correlations to future work. Within these approximations the evolution of the density operator in the interaction picture with respect to the perfect Rabi oscillation is described by
$\begin{split}\frac{{\rm d}}{{\rm d}t}{\rho}(t)=-\frac{1}{4}\int_{0}^{t}{\rm d}t^{\prime}&\{C_{\Omega}(t-t^{\prime})[\sigma_{x},\sigma_{x}\rho]\\ &+C_{\tilde{\omega}}(t-t^{\prime})[O_{z}(t^{\prime}),O_{z}(t)\rho]\}+\rm{H.c.}\end{split}$ (44)
where $C_{\Omega}(t-t^{\prime}),\text{ and }C_{\tilde{\omega}}(t-t^{\prime})$ are the auto-correlation function for each of the stochastic process and $O_{z}(t)$ is the time-dependent Pauli matrix in the frame of the ideal Hamiltonian as previously defined in Eq. (13). Projecting onto the dressed state basis {$\{\ket{+},\ket{-}\}$, we obtain again a system of equations for the diagonal terms which are decoupled from the non-diagonal terms, such that the evolution of the populations coincides exactly with the pure-dephasing case in Eq. (14).
The time evolution of the coherences can be explicitly computed within a first order Magnus expansion. Using the variables introduced in Eq. (15) the evolution reads
$\begin{split}\bm{\xi}(t)={\rm e}^{-(\Gamma_{1}(t)+\Delta\Gamma_{1}(t))}{\rm e}^{-{\rm i}\frac{\Theta(t)}{2}\bm{n}(t)\cdot\bm{\sigma}}\bm{\xi}(0),\end{split}$ (45)
where the time-dependent unit vector and angle read as before in Eqs. (16)-(17). The net effect of amplitude fluctuations amounts to a redefinition of the $T_{2}$ time of the system by adding a correction to the decay $\Gamma_{1}$ as
$\Delta\Gamma_{1}(t)=\int_{-\infty}^{\infty}\!\!\!{\rm d}\omega{S}_{\Omega}(\omega)F_{\Omega}(\omega,t),$ (46)
where we have introduced an additional decay filter function
$F_{\Omega}(\omega,t)=\frac{1}{\pi}\left(\frac{1-\cos(\omega t)}{\omega^{2}}\right)=t\eta_{\frac{2}{t}}(\omega).$ (47)
Moreover, we can generalize the closed-form expressions of the Kraus operators of the error quantum channel in Eq. (13) for the Markovian limit $\Gamma_{2}(t)=\Delta_{2}(t)=0$, to the case of multi-axis noise. The amplitude fluctuations only affect two of the Kraus operators as
$\begin{split}&K_{1}=\frac{1}{2}\sqrt{\epsilon(t)}[\sin(\Omega t)\sigma_{z}+\cos(\Omega t)\sigma_{y}]\\ &K_{2}=\frac{1}{2}\sqrt{\epsilon(t)}[\cos(\Omega t)\sigma_{z}-\sin(\Omega t)\sigma_{y}]\\ &K_{3}=\frac{1}{2}\sqrt{\xi_{-}(t)}\left[\sin\left(\frac{M_{1}(t)}{4}\right)\mathbf{1}_{2}-{\rm i}\cos\left(\frac{M_{1}(t)}{4}\right)\sigma_{x}\right]\\ &K_{4}=\frac{1}{2}\sqrt{\xi_{+}(t)}\left[\cos\left(\frac{M_{1}(t)}{4}\right)\mathbf{1}_{2}+{\rm i}\sin\left(\frac{M_{1}(t)}{4}\right)\sigma_{x}\right],\end{split}$ (48)
where the effective error rate $\epsilon(t)$ agrees with that of the pure dephasing noise from Eq. (33), and we have to introduce two additional modified error rates
$\xi_{\pm}(t)=1+{\rm e}^{-\Gamma_{1}(t)}\pm 2{\rm e}^{-\frac{1}{2}(\Gamma_{1}(t)+\Delta\Gamma_{1}(t))}.$ (49)
Figure 5: Comparison of the gate error with respect to the ideal Rabi flop for the depolarizing channel (green), the non-Clifford Kraus channel (dashed blue), and the full non-Clifford and non-Markovian channel (red) including the effects of intensity fluctuations. We also include the pure phase noise case (solid blue) to see the impact of the Rabi frequency fluctuations. The simulations use two Ornstein-Uhlenbeck processes, one for the phase noise with $\tau_{\rm c}=20\pi/\Omega$, $c=1/(40\tau_{\rm c}^{3})$, a time step $\Delta t=0.05\tau\,{\rm c}$, and a different one for the Rabi fluctuations with $\tau_{\rm c}=20\pi/\Omega$, $c=0.1/(40\tau_{\rm c}^{3})$, a time step $\Delta t=0.05\tau\,{\rm c}$, and a Rabi frequency $\Omega=2\pi\times 20\,{\rm kHz}$, which are chosen for presentation purposes.
For the non-Markovian case $\Gamma_{2}(t),\Delta_{2}(t)\neq 0$, we can again derive a closed-form analytical expression for the dynamical error map in terms of a time-dependent block-diagonal process matrix similar to Eq. (34), as can be seen in Appendix D. This analytical form reads
~$\begin{split}\tilde{\mathcal{E}}_{\rm NM}(\rho_{0})&=\sum_{\alpha,\beta=0,1}\big{[}\tilde{\chi}^{\rm err}_{A}(t)\big{]}_{\alpha\beta}E_{\alpha}\rho_{\rm id}(t)E_{\beta}^{\dagger}\\ &+\sum_{\alpha,\beta=2,3}\big{[}\tilde{\chi}^{\rm err}_{B}(t)\big{]}_{\alpha\beta}E_{\alpha}\rho_{\rm id}(t)E_{\beta}^{\dagger},\end{split}$ (50)
where the block matrices $\tilde{\chi}^{\rm err}_{A}(t),\tilde{\chi}^{\rm err}_{B}(t)$ depend on the same filtered noise PSD as the pure phase noise and, additionally, on the amplitude noise PSD via $\Delta\Gamma_{1}(t)$ (see Eqs. (120)-(121)). Again, incorporating the filters $F_{\Gamma_{2}},F_{\Delta_{2}}$ into the dynamics leads to a non-Markovian quantum dynamical map. The extent to which the map is non-Markovian can be quantified by the same measure in Eq. (35) but with a different canonical decay rate
$\bar{\gamma}_{\alpha}(t)\in\left\{0,\textstyle\frac{1}{2}\tilde{\gamma_{1}}(t)\pm\textstyle\frac{1}{2}\sqrt{\gamma_{2}^{2}(t)+\delta_{2}^{2}(t)})\right\},$ (51)
where we have replaced the decay rate $\gamma_{1}(t)$ with the new dressed decay rate $\tilde{\gamma}_{1}(t)=\gamma_{1}(t)-\delta\gamma_{1}(t)$.
Finally, we can again find the best description of the noise channel in terms of a Pauli channel by twirling. The resulting error map can be written in the form of Eq. (40) with the time-dependent error probabilities defined as
$\begin{split}p_{x}(t)&=\frac{1}{4}\left(1+{\rm e}^{-\Gamma_{1}(t)}-2\cos\textstyle\frac{1}{2}\Theta(t){\rm e}^{-\frac{1}{2}(\Gamma_{1}(t)+\Delta\Gamma_{1}(t))}\right),\\ p_{y}(t)&=\frac{1}{4}\left(1-{\rm e}^{-\Gamma_{1}(t)}+2\frac{\Delta_{2}(t)}{\Theta(t)}\sin\textstyle\frac{1}{2}\Theta(t){\rm e}^{-\frac{1}{2}{(\Gamma_{1}(t)+\Delta\Gamma_{1}(t))}}\right),\\ p_{z}(t)&=\frac{1}{4}\left(1-{\rm e}^{-\Gamma_{1}(t)}-2\frac{\Delta_{2}(t)}{\Theta(t)}\sin\textstyle\frac{1}{2}\Theta(t){\rm e}^{-\frac{1}{2}{(\Gamma_{1}(t)+\Delta\Gamma_{1}(t))}}\right),\end{split}$ (52)
and again, we have defined $p(t)=p_{x}(t)+p_{y}(t)+p_{z}(t)$.
@@ II.3 Dynamical error map with amplitude noise
II.4 Projection noise, maximum-likelihood, and analytical estimates of the gate error rate
@@
Figure 6: Reconstruction of the gate error $\overline{\epsilon}(t_{i})$ for noisy Rabi oscillations under an Ornstein-Uhlenbeck noise with $\tau_{c}=5\times 10^{-4}\,\rm{s}$ and $c=1.6\times 10^{9}\,\rm{s}^{-3}$. We include shot noise for a total number of measurements $N_{\rm shots}=2.4\cdot 10^{7}$, which allows us to approximate the tomographic probabilities $p_{s,i,\mu}$ by the corresponding relative frequencies $f_{s,i,\mu}$. The confidence region (purple shading) and the mean (purple markers) are estimated by distributing this data into $N_{r}=10^{3}$ repetitions of the maximum-likelihood reconstruction, leading to $N_{i}=1200$ shots per time step. The analytical predictions $\epsilon_{g,{\rm D}}$ in Eq. (54) (green), which is the established model in the literature, as well as the errors for our Markovian non-Clifford $\epsilon_{g,{\rm NC}}$ in Eq. (55) (blue) and non-Markovian $\epsilon_{g,{\rm NM}}$ in Eq. (56) (red) estimates, which depend on the Lorentzian noise PSD in Eq. (25). Inset: Absolute value of the difference between the ML reconstruction and the non-Clifford Kraus map (blue) or the non-Markovian map (red) infidelity.
@@
This expression captures better the oscillations of the gate error, as can be seen in the inset of Fig. 6.
Furthermore, we can give closed expressions that incorporate the effect of the amplitude fluctuations that we have discussed in the previous section. The gate infidelity in the Markovian case accounting for both phase and amplitude noise reads
$\varepsilon_{g,{\rm NC,I}}=\frac{1}{6}\left[3-{\rm e}^{-\Gamma_{1}(t)}-2{\rm e}^{-\frac{1}{2}(\Gamma_{1}(t)+\Delta\Gamma_{1}(t))}\cos\left(\textstyle\frac{1}{2}\Delta_{1}(t)\right)\right],$ (57)
which is surprisingly similar to the case of pure phase noise but with a correction to the decay rate. The non-Markovian case is entirely analogous to the purely phase noise case with the same correction such that
$\varepsilon_{g,{\rm NM,I}}=\frac{1}{6}\left[3-{\rm e}^{-\Gamma_{1}(t)}-2{\rm e}^{-\frac{1}{2}(\Gamma_{1}(t)+\Delta\Gamma_{1}(t))}\cos\textstyle\frac{1}{2}\Theta(t)\right].$ (58)
We note that the Pauli-twirled approximations $\mathcal{E}_{\rm PT}$, in Eq. (40), for both the pure dephasing noise case and including amplitude fluctuations, have the exact same fidelity as each of the non-Markovian maps respectively but, being of Pauli-type, they can be used in large-scale simulations of Clifford circuits.